# V2X-Real: a Largs-Scale Dataset for Vehicle-to-Everything Cooperative Perception

### v1

@@ 3.3 Dataset Analysis
Our dataset has an average of 36 objects per scene with a maximum of 115 and minimum of 9 objects, ensuring a high density of traffic in challenging urban scenarios. The number of annotations for each of the 10 categories is shown in Fig. 5. The dominant objects in the dataset are pedestrian and car. Compared with existing V2X dataset (Tab. 1), V2X-Real is featured with a wide variety and quantity of vulnerable road users including pedestrian, scooter, motorcycle, and bicycle. In our benchmarks, we group the 10 categories, based on their box sizes, into three super-classes i.e. pedestrian, car, and truck. In Fig. 4, we reveal the polar density map of annotated bounding boxes for these three classes, where each bin in the polar map represents a spatial area with certain distance and angle range in the ego agent’s coordinate frame. From the figure, we can observe a diverse distribution of annotations and each category is denser in the front, left, right, and rare sides of the ego vehicle. We argue this is due to the fact that most of our data is collected adjacent to the intersection and the surrounding vehicles tend to have these four relative poses with respect to the data collection vehicle. Compared with car and truck categories, pedestrian have less annotations above 100 meters due to the sparse LiDAR measurements of the small object in distant area. Fig. 6 depicts the point clouds density distribution within bounding boxes and the the bounding box size distribution. From the figure, we can see the diverse bounding box size distribution and each category will have a size distribution and could serve as a prior for object detection tasks. Additionally, the number of points within the box will decrease for longer radial distance while collaboratively viewing the objects can potentially increase the sensing observation and thus provide better visual cues.
@@ 4.1 V2X Cooperative 3D object detection
V2X-Real-I2I: In this dataset, Infrastructure-to-Infrastructure collaborations are studied, which is critical for intelligent transportation system.
Contrary to DAIR-V2X, which concentrates solely on vehicle-centric V2I collaboration, and V2V4Real, which is customized for V2V cooperative perception, the above-mentioned datasets have been meticulously crafted to cater to the wide-ranging research interests of the broader V2X community.
@@ 4.2 Metrics
Ground truth: During the training stage, a random agent will be selected as the ego agent while during the inference, the ego is fixed as per dataset categories. For example, in the V2X-Real-VC, the ego is fixed as one of the data collection vehicle while in the V2X-Real-IC, the ego is infrastructure. When the bounding boxes are not consistent across agents, we choose the ego agent’s annotation as the ground truth.
Evaluation: The evaluation is conducted in the range of -100m to 100m in x direction and -40m to 40m in y direction of the ego coordinate frame. We group the annotated 10 categories into three classes i.e. car, pedestrian and truck as per their bounding box size distribution. We calculate the Average Precision(AP) for each class based on the specified Intersection-over-Union (IoU) threshold, and a final mean Avereage Precision (mAP) is evaluated based on each category’s AP:
@@
where $C$ is the class number. As our dataset contains a substantial number of pedestrians and large buses and the bounding box sizes can vary from less than 0.5 meters (pedestrian) to over 20 meters (bus), accurately predicting object’s pose and dimension is particularly challenging. Thus similar as KITTI-360 dataset [12], we adopt smaller IoU thresholds (IoU=0.3, 0.5) for the AP calculation to faithfully reflect model’s perception capability.
Table 3: Benchmark results of SOTA cooperative perception methods for V2X-Real-VC, V2X-Real-IC, V2X-Real-V2V, and V2X-Real-I2I. All numbers in the table are represented in percentage format. For the detection accuracy, Average Precision (AP) and mean Average Precision (mAP) for car, pedestrian, and truck is reported under IoU treshold of 0.3 and 0.5.
[…]
@@ 4.3 Benchmark methods
  • No Fusion: Only ego agent’s LiDAR will be used to produce the 3D bounding boxes. It servers as the baseline.
  • Late Fusion: Each agent first generate its own predictions and then share the detection results with neighboring agents. The ego agent will receive these proposals and leverage non-maximum suppression (NMS) to produce consistent predictions.
  • Early Fusion: The raw LiDAR point clouds will be shared for agents within the communication range. After receiving neighboring agent’s LiDAR point clouds, the ego agent will transform LiDAR point clouds into ego agent’s coordinate frame and aggregate all the LiDAR point clouds to produce a holistic view for visual reasoning.
  • Intermediate Fusion: Each agent will first project its LiDAR to the ego agent’s coordinate frame and leverage LiDAR backbone to extract Birds’ Eye View (BEV) features, which are then compressed and shared with the ego agent. The ego agent will leverage received multi-agent features to refine its own feature representations. In this work, we provide benchmarks for AttFuse [26], F-Cooper [2], and V2X-ViT [25].
@@ 5.1 Implementation details
We split the dataset into train/val/test with 23379, 2770, 6850 frames respectively. For the detection models, we adopt PointPillar [7] as the LiDAR backbone. The cooperative perception framework is built upon OpenCOOD [26] and to enable multi-class cooperative perception, we extend the original single-class head to multi-class heads following OpenPCDet’s design [17]. All the models are trained with 80 epochs with a batch size of 2 and a learning rate of 0.001. Adam optimizer (decay weight $\lambda=10^{-4}$) and multi-step learning rate scheduler ($\gamma=0.1$, milestones=[10,50]) are employed for optimization.
@@ 5.2 Benchmark results
Tab. 4 summarizes the benchmark results of all six models for the four datasets. For all the dataset, cooperative methods outperform No Fusion baseline, demonstrating the great potential of V2X collaboration in enhancing the perception capabilities of individual agents. Notably, across all three object categories, the models generally perform best in the car category, whereas the performance is comparatively lower for pedestrian and truck categories. We argue this is due to the fact that truck type has a diverse dimension distribution (Fig. 5(c)) and pedestrian bounding boxes usually only contain a few sparse LiDAR points, posing challenges for the learning. Generally, intermediate fusion ranks first among three fusion strategies, and late fusion methods outperform No Fusion baselines while early fusion shows superior performance than late fusion.
@@ 6 Conclusion
In this appendix, we first present additional visualizations for V2X-Real dataset (Sec. 0.A). Then, we detail the model configurations for LiDAR-based and Camera-based Cooperative Perception methods (Sec. 0.B). Afterwords, additional benchmark results and qualitative visualizations are shown in Sec. 0.C.
Appendix 0.A Dataset Visualization
V2X-Real is collected in two types of scenarios i.e. V2X smart intersection and V2V corridors. Visualization of smart intersection are presented in the Fig. 1-3 of the main paper. In this appendix, additional visualizations are provided for the V2V corridors (Fig. 7). Each figure displays three images: the top image showcases the aggregated LiDAR point clouds from both vehicles, while the bottom images present views from the associated front-view cameras. Further, Fig. 8 illustrates the stereo image pairs from the front camera. Notably, each vehicle is equipped with four stereo cameras, yielding eight images per frame, as illustrated in Fig. 9. These cameras, each offering a 120-degree horizontal field of view (FoV), are mounted in four perpendicular directions to achieve a complete 360-degree FoV around the vehicle. In Fig. 9, ground-truth 3D bounding boxes are projected onto the image plane via the intrinsics and extrinsics and each category i.e., pedestrian, car, and truck, are colored differently.
@@ Appendix 0.A Dataset Visualization
Figure 9: Visualization of four stereo cameras of the vehicle. (a)-(d) plot the images from front, right, rear, and left camera respectively. 3D bounding boxes are projected onto 2D image plane and are colored differently as per category. Best viewed in color with zoom for enhanced detail.
@@ 0.B.1 LiDAR-based Cooperative Perception
LiDAR backbone: We adopt PointPillar [7] as the LiDAR backbone with a voxel size of 0.4 meters in x and y directions. Following [25, 26, 20], each agent first project its LiDAR to the ego agent’s coordinate frame. Then the raw points are converted into pillar features and scattered into a 2D Bird’s Eye View (BEV) feature map, which is shared with neighboring connected agents. In this work, the number of points per voxel is set to 32 while the maximum voxel number is 64000.
Fusion method: The ego agent receives all the collaborator’s feature and produce refined feature representation via fusing all the received feature maps and its own extracted features. We provide benchmarks for three intermediate fusion methods i.e., F-Cooper [2], AttFuse [26], V2X-ViT [25]. F-Cooper leverages max-pooling to aggregate features. AttFuse adopts per-location attention for feature refinement. V2X-ViT employs a multi-scale window attention and heterogeneous self-attention to jointly learn the inter-agent and intra-agent interactions.
Head: We adopt two 1x1 convolution layers for box regression and classification. The regression branch outputs $(x,y,z,w,h,l,\theta)$, denoting the position $(x,y,z)$, box dimensions $(w,h,l)$ and yaw angle $(\theta)$ with respect to the predefined anchor box of each category. For the classification branch, the head outputs the confidence score of each category i.e., pedestrian, car, truck. To capture the dimension prior of various types of objects, we define a set of fixed-size anchor box for each category as per its box size distribution and the network only learns the displacement with respect to these predefined anchor boxes. In this way, the model can efficiently identify multi-scale objects (Fig. 10).
@@ 0.B.2 Camera-based Cooperative Perception
For saving the computation, only the left camera from each stereo camera pair is used, leading to a total of 4 images (Fig. 9) per vehicle. To extract salient features from 2D images, we employ a similar architecture as BEVFormer [11] with no temporal information. Following [20], for faster running speed, we adopt ResNet50 to extract 2D image features and leverage anchor-based head to generate the final detection outputs. In camera-based cooperative perception, we provide benchmarks for No Fusion and Late Fusion methods.
@@
The evaluation for the camera perception is conducted in the range of 50 meters in x and y directions of ego coordinate frame.
Table 4: Additional benchmark results of SOTA cooperative perception methods for V2X-Real-V2V and V2X-Real-I2I. All numbers in the table are represented in percentage format.
[…]
0.3 0.5 0.3 0.5 0.3 0.5 0.3 0.5
V2X-Real-V2V No Fusion 41.7 39.4 26.9 14.4 21.6 13.7 30.1 22.5
Late Fusion 47.4 44.4 29.2 14.9 18.7 9.1 31.8 22.8
Early Fusion 54.0 49.8 31.9 17.1 28.6 18.6 38.1 28.5
F-Cooper [2] 42.7 40.3 27.7 14.0 25.6 18.6 32.0 24.3
AttFuse [26] 58.6 55.3 30.1 15.4 28.9 21.7 39.2 30.8
V2X-ViT [25] 59.0 56.3 37.4 20.7 42.9 35.0 46.5 37.3
V2X-Real-I2I No Fusion 48.6 40.0 30.9 15.8 23.5 22.4 34.3 26.1
Late Fusion 67.2 63.3 41.1 23.1 48.4 39.1 52.2 41.8
Early Fusion 60.9 57.2 41.2 21.9 38.5 30.8 46.9 36.6
F-Cooper [2] 71.5 65.5 49.3 27.5 50.0 40.7 56.9 44.6
AttFuse [26] 73.5 69.5 42.8 20.5 53.2 39.4 56.5 43.1
V2X-ViT [25] 77.3 68.8 54.4 30.5 56.2 51.5 62.7 50.3
@@ 0.C.1 LiDAR-based Cooperative Perception
Tab. 4 shows the additional benchmark results for V2X-Real-V2V and V2X-Real-I2I. All the fusion method outperform No Fusion baselines. Among these, Intermediate Fusion methods typically yield more precise predictions compared to Late and Early Fusion strategies. Fig. 10 offers qualitative visualizations for all the benchmarked methods, confirming the overarching trend observed in the main paper—fusion methods excel over the No Fusion baseline, with Intermediate Fusion methods showing superior precision. Notably, while Late Fusion successfully identifies more true positives compared with No Fusion, it also leads to a rise in false negative predictions. We argue this is due to the loss of context during the late fusion strategy where only the detection results are shared and fused. Consequently, the uncertainty inherent in collaborator messages may be amplified, resulting in an increased rate of false negatives. On the other hand, intermediate fusion methods produce much less false positives. We contend that this is due to the end-to-end multi-agent fusion, where the shared/fused features preserve the raw context information and thus can better help refine the ego agent’s feature representation, leading to boosted performance.
@@ 0.C.2 Camera-based Cooperative Perception
In this work, we provide benchmarks for vehicle-side multi-view camera-based cooperative perception and leave infrastructure-side camera solution for future works as infrastructure camera has fixed topology with roads, which could serve as a prior for detection task and thus require a dedicated new algorithm for efficient Infrastructure-to-Everything cooperative perception. Tab. 5 demonstrates the benchmark performance for camera-based cooperative perception on V2X-Real-VC (excluding infrastructure collaborator) and V2X-Real-V2V. The results indicate that Late Fusion outperforms No Fusion baseline for both datasets. Moreover, both models perform higher in the V2X-Real-VC dataset than the V2X-Real-V2V, which contains V2V corridor scenarios. We argue this is due to the fact that the V2V corridors contain more diverse scenarios and contain roads with high slope changes, posing challenges for BEV-based camera solutions in accurately determining 3D poses. We encourage further exploration into these important yet under-explored areas by the research community.

### v2

@@ 3.3 Dataset Analysis
Our dataset has an average of 36 objects per scene with a maximum of 115 and a minimum of 9 objects, ensuring a high density of traffic in challenging urban scenarios. The number of annotations for each of the 10 categories is shown in Fig. 5. The dominant objects in the dataset are pedestrians and cars. Compared with the existing V2X dataset (Tab. 1), V2X-Real is featured with a wide variety and quantity of vulnerable road users including pedestrian, scooter, motorcycle, and bicycle. In our benchmarks, we group the 10 categories, based on their box sizes, into three super-classes i.e. pedestrian, car, and truck. In Fig. 4, we reveal the polar density map of annotated bounding boxes for these three classes, where each bin in the polar map represents a spatial area with a certain distance and angle range in the ego agent’s coordinate frame. From the figure, we can observe a diverse distribution of annotations and each category is denser in the front, left, right, and rare sides of the ego vehicle. We argue this is due to the fact that most of our data is collected adjacent to the intersection and the surrounding vehicles tend to have these four relative poses with respect to the data collection vehicle. Compared with car and truck categories, pedestrians have fewer annotations above 100 meters due to the sparse LiDAR measurements of the small objects in distant areas. Fig. 6 depicts the point clouds density distribution within bounding boxes and the bounding box size distribution. From the figure, we can see the diverse bounding box size distribution and each category will have a size distribution and could serve as a prior for object detection tasks. Additionally, the number of points within the box will decrease for longer radial distances while collaboratively viewing the objects can potentially increase the sensing observation and thus provide better visual cues.
@@ 4.1 V2X Cooperative 3D object detection
V2X-Real-I2I: In this dataset, Infrastructure-to-Infrastructure collaborations are studied, which is critical for intelligent transportation systems.
Contrary to existing works tailored for a single collaboration mode, the presented datasets have been meticulously crafted to cater to the diverse research interests of the broader V2X community.
@@ 4.2 Metrics
Ground truth: During the training stage, a random agent will be selected as the ego agent while during the inference, the ego is fixed as per dataset categories. For example, in the V2X-Real-VC, the ego is fixed as one of the data collection vehicles while in the V2X-Real-IC, the ego is infrastructure. When the bounding boxes are not consistent across agents, we choose the ego agent’s annotation as the ground truth.
Evaluation: The evaluation is conducted in the range of -100m to 100m in x direction and -40m to 40m in y direction of the ego coordinate frame. We group the annotated 10 categories into three classes i.e. car, pedestrian and truck as per their bounding box size distribution. We calculate the Average Precision(AP) for each class based on the specified Intersection-over-Union (IoU) threshold, and a final mean Average Precision (mAP) is evaluated based on each category’s AP:
@@
where $C$ is the class number. As our dataset contains a substantial number of pedestrians and large buses and the bounding box sizes can vary from less than 0.5 meters (pedestrian) to over 20 meters (bus), accurately predicting objects’ poses and dimensions is particularly challenging. Thus similar as KITTI-360 dataset [15], we adopt smaller IoU thresholds (IoU=0.3, 0.5) for the AP calculation to faithfully reflect model’s perception capability.
Table 3: Benchmark results of SOTA cooperative perception methods for V2X-Real-VC, V2X-Real-IC, V2X-Real-V2V, and V2X-Real-I2I. All numbers in the table are represented in percentage format. For detection accuracy, Average Precision (AP) and mean Average Precision (mAP) for cars, pedestrians, and trucks are reported under IoU thresholds of 0.3 and 0.5.
Dataset Models APcar@IoU APped.@IoU APtruck@IoU mAP@IoU
@@ 4.3 Benchmark methods
  • No Fusion: Only ego agent’s LiDAR will be used to produce the 3D bounding boxes. It serves as the baseline.
  • Late Fusion: Each agent first generates its own predictions and then shares the detection results with neighboring agents. The ego agent will receive these proposals and leverage non-maximum suppression (NMS) to produce consistent predictions.
  • Early Fusion: The raw LiDAR point clouds will be shared for agents within the communication range. After receiving the neighboring agent’s LiDAR point clouds, the ego agent will transform LiDAR point clouds into the ego agent’s coordinate frame and aggregate all the LiDAR point clouds to produce a holistic view for visual reasoning.
  • Intermediate Fusion: Each agent will first project its LiDAR to the ego agent’s coordinate frame and leverage the LiDAR backbone to extract Birds’ Eye View (BEV) features, which are then compressed and shared with the ego agent. The ego agent will leverage received multi-agent features to refine its own feature representations. In this work, we provide benchmarks for AttFuse [35], F-Cooper [2], and V2X-ViT [34].
@@ 5.1 Implementation details
We split the dataset into train/val/test with 23379, 2770, and 6850 frames respectively. For the detection models, we adopt PointPillar [9] as the LiDAR backbone. The cooperative perception framework is built upon OpenCOOD [35]. To enable multi-class cooperative perception, we extend the original single-class head to multi-class heads following OpenPCDet’s design [21]. All the models are trained with 80 epochs with a batch size of 2 and a learning rate of 0.001. Adam optimizer (decay weight $\lambda=10^{-4}$) and multi-step learning rate scheduler ($\gamma=0.1$, milestones=[10,50]) are employed for optimization.
@@ 5.2 Benchmark results
Tab. 3 summarizes the benchmark results of all six models for the four datasets. For all the datasets, cooperative methods outperform No Fusion baseline, demonstrating the great potential of V2X collaboration in enhancing the perception capabilities of individual agents. Among these cooperative methods, Intermediate Fusion methods typically yield more precise predictions compared to Late and Early Fusion strategies. Notably, across all three object categories, the models generally perform best in the car category, whereas the performance is comparatively lower for the pedestrian and truck categories. We argue this is due to the fact that truck type has a diverse dimension distribution (Fig. 6(c)) and pedestrian bounding boxes usually only contain a few sparse LiDAR points, posing challenges for the learning. Generally, intermediate fusion ranks first among three fusion strategies, and late fusion methods outperform No Fusion baselines while early fusion shows superior performance than late fusion.
@@ 6 Conclusion
Future work and limitation: The current dataset comprises only LiDAR and camera data, lacking support for radar, which is crucial for developing low-cost, scalable autonomous driving solutions. In the future, we plan to further extend our sensor suits to low-cost sensors such as radar and low-beam LiDAR for studying scalable low-cost V2X solutions.
Acknowledgements
Federal Highway Administration Center of Excellence on New Mobility and Automated Vehicles, National Science Foundation # 2346267 POSE: Phase II: DriveX: An Open-Source Ecosystem for Automated Driving and Intelligent Transportation Research
@@ Acknowledgements
In this appendix, we first present additional visualizations for the V2X-Real dataset (Sec. 0.A). Then, we detail the model configurations for LiDAR-based and Camera-based Cooperative Perception methods (Sec. 0.B). Afterwords, additional benchmark results, and qualitative visualizations are shown in Sec. 0.C.
@@
Appendix 0.A Dataset Visualization
V2X-Real is collected in two types of scenarios i.e. V2X smart intersection and V2V corridors. Visualizations of the smart intersection are presented in Fig. 1-3 of the main paper. In this supplementary material, additional visualizations are provided for the V2V corridors (Fig. 7). Each figure displays three images: the top image showcases the aggregated LiDAR point clouds from both vehicles, while the bottom images present views from the associated front-view cameras. Further, Fig. 8 illustrates the stereo image pairs from the front camera. Notably, each vehicle is equipped with four stereo cameras, yielding eight images per frame, as illustrated in Fig. 9. These cameras, each offering a 120-degree horizontal field of view (FoV), are mounted in four perpendicular directions to achieve a complete 360-degree FoV around the vehicle. In Fig. 9, ground-truth 3D bounding boxes are projected onto the image plane via the intrinsics and extrinsics and each category i.e., pedestrian, car, and truck, are colored differently.
@@ Appendix 0.A Dataset Visualization
Figure 9: Visualization of four stereo cameras of the vehicle. (a)-(d) plot the images from the front, right, rear, and left cameras respectively. 3D bounding boxes are projected onto a 2D image plane and are colored differently as per category. Best viewed in color with zoom for enhanced detail.
@@ 0.B.1 LiDAR-based Cooperative Perception
LiDAR backbone: We adopt PointPillar [9] as the LiDAR backbone with a voxel size of 0.4 meters in the x and y directions. Following [34, 35, 26], each agent first project its LiDAR to the ego agent’s coordinate frame. Then, the raw points are converted into pillar features and scattered into a 2D Bird’s Eye View (BEV) feature map, which is shared with neighboring connected agents. In this work, the number of points per voxel is set to 32 while the maximum voxel number is 64000.
Fusion method: The ego agent receives all the collaborators’ features and produces refined feature representation via fusing all the received feature maps and its own extracted features. We provide benchmarks for three intermediate fusion methods i.e., F-Cooper [2], AttFuse [35], V2X-ViT [34]. F-Cooper leverages max-pooling to aggregate features. AttFuse adopts per-location attention for feature refinement. V2X-ViT employs multi-scale window attention and heterogeneous self-attention to jointly learn the inter-agent and intra-agent interactions.
Head: We adopt two 1x1 convolution layers for box regression and classification. The regression branch outputs $(x,y,z,w,h,l,\theta)$, denoting the position $(x,y,z)$, box dimensions $(w,h,l)$ and yaw angle $(\theta)$ with respect to the predefined anchor box of each category. For the classification branch, the head outputs the confidence score of each category i.e., pedestrian, car, truck. To capture the dimension prior of various types of objects, we define a set of fixed-size anchor boxes for each category as per its box size distribution and the network only learns the displacement with respect to these predefined anchor boxes. In this way, the model can efficiently identify multi-scale objects (Fig. 10).
@@ 0.B.2 Camera-based Cooperative Perception
For saving the computation, only the left camera from each stereo camera pair is used, leading to a total of 4 images (Fig. 9) per vehicle. To extract salient features from 2D images, we employ an architecture similar to BEVFormer [14] with no temporal information. Following [26], for faster running speed, we adopt ResNet50 to extract 2D image features and leverage anchor-based head to generate the final detection outputs. In camera-based cooperative perception, we provide benchmarks for No Fusion and Late Fusion methods.
@@
The evaluation for the camera perception is conducted in the range of 50 meters in the x and y directions of the ego coordinate frame.
@@ 0.C.1 LiDAR-based Cooperative Perception
Fig. 10 offers qualitative visualizations for all the benchmarked methods, confirming the overarching trend observed in the main paper—fusion methods excel over the No Fusion baseline, with Intermediate Fusion methods showing superior precision. Notably, while Late Fusion successfully identifies more true positives compared with No Fusion, it also leads to a rise in false negative predictions. We argue this is due to the loss of context during the late fusion strategy where only the detection results are shared and fused. Consequently, the uncertainty inherent in collaborator messages may be amplified, resulting in an increased rate of false negatives. On the other hand, intermediate fusion methods produce much fewer false positives. We contend that this is due to the end-to-end multi-agent fusion, where the shared/fused features preserve the raw context information and thus can better help refine the ego agent’s feature representation, leading to boosted performance.
@@ 0.C.2 Camera-based Cooperative Perception
In this work, we provide benchmarks for vehicle-side multi-view camera-based cooperative perception and leave infrastructure-side camera solution for future works as the infrastructure camera has a fixed topology with roads, which could serve as a prior for detection task and thus require a dedicated new algorithm for efficient Infrastructure-to-Everything cooperative perception. Tab. 4 demonstrates the benchmark performance for camera-based cooperative perception on V2X-Real-VC (excluding infrastructure collaborator) and V2X-Real-V2V. The results indicate that Late Fusion outperforms No Fusion baseline for both datasets. Moreover, both models perform higher in the V2X-Real-VC dataset than the V2X-Real-V2V, which contains V2V corridor scenarios. We argue this is due to the fact that the V2V corridors contain more diverse scenarios and roads with high slope changes, posing challenges for BEV-based camera solutions in accurately determining 3D poses. We encourage further exploration into these important yet under-explored areas by the research community.